A pendulum close to the surface of the earth oscillates with a period of 3.76 seconds. How long is the pendulum?
1
Expert's answer
2020-10-12T07:48:54-0400

We can find the length of the pendulum from the formula:


"T = 2 \pi \sqrt{\dfrac{L}{g}},"

here, "T = 3.76 \ s" is the period of the pendulum. "L" is the length of the pendulum and "g = 9.8 \ \dfrac{m}{s^2}" is the acceleration due to gravity.

Then, from this formula we can find the length of the pendulum:


"L = \dfrac{T^2 g}{4 \pi^2} = \dfrac{(3.76 \ s)^2 \cdot 9.8 \ \dfrac{m}{s^2}}{4 \pi^2} = 3.51 \ m."

Answer:

"L = 3.51 \ m."